Part A Find the current in each resistor in the fi
Answer #1

Here ,

let the current in R1 is I1

current in R2 is I2

and current in R3 is I3

Using KCL at A

I1 + I2 - I3 = 0 ----(1)

Using KVL in the left loop ,

9 - 4 I1 - 2.6 I3 - 9 I1= 0

13 I1 + 2.6 I3 = 9 ----(2)

using KVL in the right loop ,

12 - 5.2 I2 - 2.6 I3 = 0

5.2 I2 + 2.6 I3 = 12 ----(3)

solving 1 , 2 , 3

I1 = 0.34 A

I2 = 1.43 A

I3 = 1.76 A

as I4 = I1

I4 = 0.34 A

Hence ,

currents are (0.34 A , 1.43 A , 1.76 A , 0.34 A)